OpenStudy (anonymous):

Find sin(x) if cos(x)=a/a+2.

OpenStudy (anonymous):

a) \[\frac{ 2\sqrt{a+1} }{a+2 }\] b) \[\frac{ 4(a+1) }{a+2 }\] c) \[\frac{ 2\sqrt{a+1} }{a }\] d) \[\frac{ 4(a+1) }{ a }\]

OpenStudy (anonymous):

e) \[\frac{ \sqrt{a+1} }{ 2a }\]

OpenStudy (mertsj):

\[\sin ^2x+\cos ^2x=1\]

OpenStudy (mertsj):

\[\sin ^2x+(\frac{a}{a+2})^2=1\]

OpenStudy (mertsj):

\[\sin ^2x=1-\frac{a^2}{(a+2)^2}=\frac{(a+2)^2-a^2}{(a+2)^2}=\]

OpenStudy (mertsj):

\[\frac{a^2+4a+4-a^2}{(a+2)^2}=\frac{4(a+1)}{(a+2)^2}\]

OpenStudy (mertsj):

\[\sin x=\frac{2\sqrt{a+1}}{a+2}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (mertsj):

yw